# PyEncode: An Open-Source Library for Structured Quantum State Preparation

### 30 Mar 2026 v1

@@ 2 Prior Work
The works cited above provide theoretical constructions without open-source implementations. PyEncode fills this gap, providing a unified, immediately deployable library that implements these constructions directly: SPARSE via Gleinig–Hoefler [8]; STEP via Shukla–Vedula [24]; WALSH via Welch et al. [27]; and FOURIER via the exact, finite-term inverse-QFT pipeline of Gonzalez-Conde et al. and Moosa et al. [9, 20]. The SQUARE synthesizer extends STEP to general intervals $k_{1},k_{2})$ by composing a Shukla–Vedula step circuit with a Draper QFT-based constant adder [6], at $\mathcal{O}(m^{2})$ cost in general and $\mathcal{O}(m)$ for power-of-2-aligned blocks. The GEOMETRIC synthesizer packages the well-known product-state structure of $r^{i}$ [28] as a named pattern, preparing exponential-decay sequences in $\mathcal{O}(m)$ single-qubit gates with zero entangling gates. The LCU constructor implements Protocol 1 LCU [3] for weighted superpositions of pattern states, with analytically determined success probability.
@@ 3.4 Fourier
Example. A pure sine wave is a single Fourier mode with $\varphi=0$. The DFT has two nonzero coefficients; PyEncode prepares that sparse Fourier state and applies the inverse QFT:
@@ 3.5 Walsh
# Generalized: two distinct levels, no ancilla
@@ 3.6.1 Construction
3.7 Linear Combination of Unitaries
@@ 5 Gate Count Comparison
Figure 10: Transpiled gate count vs. number of qubits $m$ ($N=2^{m}$), for $m=4,6,8,10,12$. All PyEncode patterns remain below 30 gates across the full range, while Qiskit StatePreparation on a random vector grows from 26 gates at $m=4$ to 8,178 at $m=12$. SPARSE uses $s=2$ entries with non-aligned indices; SQUARE uses a non-aligned interval [N/4+1, 3N/4+1)[N/4{+}1,\;3N/4{+}1), which activates the general $\mathcal{O}(m^{2})$ Draper-adder path. All patterns scale as $\mathcal{O}(\mathrm{poly}(m))$.
@@ 6.1 Quantum Chemistry
Fermi-Hubbard model. After a Jordan-Wigner transformation on an $L$-site chain [15, 16], under a suitable term ordering, the Pauli coefficient vector takes exactly two distinct values: the hopping amplitude $t$ on the first $L$ terms and the on-site interaction $U$ on the remaining $L$ terms ($N=2L$ total). This two-valued equal-block structure is exactly a generalized Walsh function, giving an ancilla-free single-register PREP circuit in $m+1=\log_{2}(2L)+1$ gates:
@@
Figure 11 shows the coefficient vector and circuit.
@@
In both cases the gate saving over Qiskit’s general preparation is exponential, and it compounds over the many PREP–SELECT repetitions required by phase estimation [3].
@@ 6.2 Computational Mechanics
$\mathbf{f}=\mathbf{u}\otimes\mathbf{v},\qquad u_{i}=\sin(2\pi ni/N),\quad v_{j}=\sin(2\pi pj/N).$ (13)
@@ 6.3 Quantum Finance
Quantum amplitude estimation (QAE) offers a quadratic speedup over classical Monte Carlo for computing expectations $\mathbb{E}[f(X)]$ [19]. A common special case is a discrete approximation to the price distribution with $s$ mass points, handled directly by the SPARSE constructor:
@@
SPARSE(list(zip(indices, weights))),
N=N)
# info.complexity -> "O(s*m)"
Figure 13 shows the discretized price distribution.
Figure 13: Left: Price distribution as SPARSE (s=3s\!=\!3, N=8N\!=\!8) Right: PyEncode circuit.
Since QAE applies the state preparation oracle $\mathcal{O}(1/\varepsilon)$ times [19], a reduction from $\mathcal{O}(2^{m})$ to $\mathcal{O}(|S|m)$ substantially reduces total circuit depth.
@@ 7 Conclusions
  • LCU — Protocol 1 LCU [3] for weighted superpositions of $r$ component states, with analytically computed success probability and automatic disjoint-support detection.

### 02 Apr 2026 v3

@@ 2 Prior Work
The works cited above provide theoretical constructions without open-source implementations. PyEncode fills this gap, providing a unified, immediately deployable library that implements these constructions directly: SPARSE via Gleinig–Hoefler [8]; STEP via Shukla–Vedula [24]; WALSH via Welch et al. [27]; and FOURIER via the exact, finite-term inverse-QFT pipeline of Gonzalez-Conde et al. and Moosa et al. [9, 20]. The SQUARE synthesizer extends STEP to general intervals $k_{1},k_{2})$ by composing a Shukla–Vedula step circuit with a Draper QFT-based constant adder [6], at $\mathcal{O}(m^{2})$ cost in general and $\mathcal{O}(m)$ for power-of-2-aligned blocks. The GEOMETRIC synthesizer packages the well-known product-state structure of $r^{i}$ [28] as a named pattern, preparing exponential-decay sequences in $\mathcal{O}(m)$ single-qubit gates with zero entangling gates. The LCU constructor implements LCU [3] for weighted superpositions of pattern states, with analytically determined success probability.
@@ 3.4 Fourier
Example. A pure sine wave is a single Fourier mode with $\varphi=0$. The DFT has two nonzero coefficients; PyEncode prepares a sparse Fourier state and applies the inverse QFT:
@@ 3.5 Walsh
# Two distinct levels, no ancilla
@@ 3.6.1 Construction
3.7 LCU (Linear Combination of Unitaries)
@@ 5 Gate Count Comparison
Figure 10: Transpiled gate count vs. number of qubits $m$ ($N=2^{m}$), for $m=4,6,8,10,12$. All PyEncode patterns remain below 30 gates across the full range, while Qiskit StatePreparation on a random vector grows from 26 gates at $m=4$ to 8,178 at $m=12$. SPARSE uses $s=2$ entries with non-aligned indices; SQUARE uses a non-aligned interval [N/4+1, 3N/4+1)[N/4{+}1,\;3N/4{+}1), which activates the general $\mathcal{O}(m^{2})$ Draper-adder path. All patterns scale as $\mathcal{O}(\mathrm{poly}(m))$. Circuit depth follows the same asymptotic scaling; see Table 2 for per-pattern depth at $m=6$.
@@ 6.1 Quantum Chemistry
Fermi-Hubbard model. After a Jordan-Wigner transformation on an $L$-site chain [15, 16], the Pauli coefficient vector takes exactly two distinct values: the hopping amplitude $t$ on the first $L$ terms and the on-site interaction $U$ on the remaining $L$ terms ($N=2L$ total). This two-valued equal-block structure is exactly a generalized Walsh function, giving an ancilla-free single-register PREP circuit in $m+1=\log_{2}(2L)+1$ gates:
@@
Figure 11 shows the coefficient vector and circuit. Although the coefficient vector resembles two concatenated step functions, the generalized Walsh synthesizer prepares it in a single $m+1$-gate circuit without ancilla or post-selection.
@@
In both cases, the gate saving over Qiskit’s general preparation is exponential, and it applies to every PREP–SELECT repetition required by phase estimation [3].
In passing, we note that the Pauli coefficient vector of the 1D Poisson stiffness matrix also admits efficient encoding: its magnitudes follow a geometric series $2^{-k}$, directly handled by the GEOMETRIC pattern.
@@ 6.2 Computational Mechanics
$\mathbf{f}=\mathbf{u}\otimes\mathbf{v},u_{i}=\sin(2\pi ni/N),v_{j}=\sin(2\pi pj/N).$ (13)
@@ 6.3 Quantum Finance
Quantum amplitude estimation (QAE) offers a quadratic speedup over classical Monte Carlo for computing expectations $\mathbb{E}[f(X)]$ [19]. The state preparation oracle must load the discretized probability distribution into a quantum register. A piecewise-constant approximation — where the price axis is divided into bins of equal width, each with a distinct probability mass — is a natural fit for the LCU constructor with disjoint SQUARE components:
@@
LCU([(0.10, SQUARE(k1=0, k2=4, c=1.0)),
(0.40, SQUARE(k1=4, k2=8, c=1.0)),
(0.35, SQUARE(k1=8, k2=12, c=1.0)),
(0.15, SQUARE(k1=12, k2=16, c=1.0))]),
N=16)
# info.success_probability -> 0.315
PyEncode detects that the four intervals are disjoint and computes $p=0.315$ analytically via (11). Figure 13 shows the 4-bin distribution and the LCU circuit.
Figure 13: Left: Price distribution as LCU of 4 disjoint SQUARE bins, N=16N=16Right: PyEncode circuit.
Since QAE applies the state preparation oracle $\mathcal{O}(1/\varepsilon)$ times [19], the $\mathcal{O}(\mathrm{poly}(m))$ circuit cost of each LCU component compounds into substantial savings over a general $\mathcal{O}(2^{m})$ preparation.
@@ 7 Conclusions
  • LCU — LCU [3] for weighted superpositions of $r$ component states, with analytically computed success probability and automatic disjoint-support detection.
@@ Acknowledgments
Use of AI Tools
The authors used generative AI assistants including Claude (Anthropic) and Gemini (Google) during the preparation of this manuscript. The tools were used for code development and drafting of text. All content was reviewed, verified, and edited by the authors, who take full responsibility for the accuracy and integrity of the work.